OpenStudy (anonymous):

Veronica Foaley sells evening jackets sewn with metallic threads and beads. She wants to set a price of $600 for a jacket. The markup she wants is 60% of the selling price. What is the most she can pay for the jacket?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@andrewhaze

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

@justine0426

OpenStudy (anonymous):

@ChipperJay*

OpenStudy (anonymous):

@alexalas

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

If I call with p the selling price, then we have: \[\frac{{60}}{{100}} \times p = 600\]

OpenStudy (anonymous):

mhmm

OpenStudy (michele_laino):

is $600 the selling price, right?

OpenStudy (anonymous):

yeah

OpenStudy (michele_laino):

ok! then the 60% of the selling price is: \[\frac{{60}}{{100}} \times 600 = ...?\]

OpenStudy (anonymous):

360

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

it says its wrong

OpenStudy (michele_laino):

Please wait, I think that your answer is: 600-360=240

OpenStudy (michele_laino):

is it right!

OpenStudy (michele_laino):

sorry: is it right?

OpenStudy (anonymous):

yes!!! haha :)

OpenStudy (anonymous):

A store owner wants a markup of 34% on the selling price of goods. If the selling price is $1.50 for each cookie, what is the most the store owner should pay per dozen cookies?

OpenStudy (anonymous):

canu help mew with this please

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

here the selling price for a dozen of cookies, is: 1.50*12=...?

OpenStudy (anonymous):

18

OpenStudy (michele_laino):

ok! now the 34% of the selling price, is: \[\frac{{34}}{{100}} \times 18 = ...?\]

OpenStudy (anonymous):

6.12

OpenStudy (michele_laino):

that's right! so your answer is: 18-6.12=...?

OpenStudy (anonymous):

11.;88

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

sJananda Roti owns a flower shop and marks items up at least 35% on the selling price. What is the most Jananda can pay for a dozen roses, if she sells each rose for $2.00?

OpenStudy (anonymous):

do i 35/100x 2.00

OpenStudy (michele_laino):

we have to multiply 2 by 12, namely 2*12=...?

OpenStudy (anonymous):

24

OpenStudy (michele_laino):

ok! now you can compute this: 35/100*24=...?

OpenStudy (anonymous):

8.4

OpenStudy (michele_laino):

ok! the requested answer is: 24-8.4=...?

OpenStudy (anonymous):

15.6

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

omg thank you so much you have been a great help and helped clarify

OpenStudy (michele_laino):

thank you!